Rockfalls and small pyroclastic flows continue to be generated from the new dome which is growing very quickly on top of the October 1 dome. Seismicity is at a variable level, with irregular swarms of hybrid events and rockfall signals dominating the records. Progressively larger pyroclastic flows are anticipated in the Tar River valley, and large collapses of the new dome and older dome material could occur quite suddenly. Ash generated from pyroclastic flows will be blown by the wind, probably towards the west. The instability of the Galway's Wall remains of concern, and the possibility of rapid escalation to explosive activity still exists.
There is no access to zone A/B. Restricted access is allowed to zone C/D, but only for essential purposes and by people with a means of rapid exit. All other zones have normal occupation.
Observations were limited during the day today due to low cloud. The observations made do suggest that rockfall activity is occurring on a semi-continuous basis although run out is limited to the base of the dome for all but the largest rockfalls. Excellent viewing conditions last evening showed the very active nature of the new dome, with glowing seen all over the northeastern flank and common avalanching of incandescent blocks. The new dome is very dark in colour and is quite smooth compared to older domes - a sample is required before detailed analysis of the rock can be undertaken.
Seismic activity comprises mainly hybrid events in poorly-defined swarms at intervals of a few hours. There were 138 hybrid events recorded during the day, along with 6 long-period and two volcano-tectonic earthquakes. A total of 36 rockfalls triggered the seismic network today, a significant reduction from yesterday. Although the number of events was lower today than yesterday, individual events have tended to be bigger.
No EDM measurements were taken today due to the low cloud cover.
Water and gas samples were collected today as part of the routine environmental monitoring undertaken by MVO.
COSPEC measurements made yesterday were further processed overnight and an average flux of 325 tonnes per day of sulphur dioxide was measured. This is similar to the previous day and also similar to measurements in mid-November when COSPEC was last operating.
Temperature measurements on the new dome were attempted yesterday evening but the instrument used needed to be closer than was possible in the conditions. However, observations of incandescence in daylight suggest that this dome lava may be hotter than previous dome lava.